# Towards a generic absoluteness theorem for Chang models

### v2

@@ Abstract.
Inspired by core model induction, we introduce the definable powerset $\mathcal{A}^{\infty}$ of $\Gamma^{\infty}$ and use our derived model representation mentioned above to show that the theory of $L(\mathcal{A}^{\infty})$ cannot be changed by forcing (see Theorem 1.19). Working in a different direction, we also show that the theory of $L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}]$, where $\mathcal{C}$ is the club filter on ${\wp}_{\omega_{1}}(\Gamma^{\infty})$, cannot be changed by forcing (see Theorem 1.30). Proving the two aforementioned results is the first step towards showing that the theory of $L(Ord^{\omega},\Gamma^{\infty},\mathbb{R})([\mu_{\alpha}:\alpha\in Ord])$, where $\mu_{\alpha}$ is the club filter on ${\wp}_{\omega_{1}}(\alpha)$, cannot be changed by forcing.
The first author gratefully acknowledges the support of the Elise Richter grant number V844, START grant number Y1498, and the international grant number I6087 of the Austrian Science Fund (FWF). The second author’s work is funded by the National Science Centre, Poland under the Weave-UNISONO call in the Weave program, registration number UMO-2021/03/Y/ST1/00281.
@@ 1.1. Main results
Following Woodin, we write $\Gamma^{\infty}$ for the set of universally Baire sets and, if $g$ is set generic over $V$, we write $\Gamma_{g}^{\infty}=(\Gamma^{\infty})^{V[g]}$ and $\mathbb{R}_{g}=\mathbb{R}^{V[g]}$ where $\mathbb{R}$ denotes the set of reals. Given a universally Baire set $A$ and a set generic $g\subseteq\mathbb{P}$ for some poset $\mathbb{P}$, we write $A_{g}$ for the interpretation of $A$ in $V[g]$. More precisely, letting $\kappa\geq{|\mathbb{P}|}^{+}$ be a cardinal and $(S,T)$ be any $\kappa$-absolutely complementing trees with $p[S]=A$, we set $A_{g}=(p[S])^{V[g]}$. It can be easily checked using the absoluteness of well-foundedness that $A_{g}$ is independent of $(T,S)$. The main object this paper studies is the model $L(\Gamma^{\infty},\mathbb{R})$, and our aim is to find a useful derived model representation for it [^1] By a theorem of Woodin, all models of ${\sf{AD^{+}}}$ are elementary equivalent to a model of ${\sf{AD^{+}}}$ that can be obtained as a derived model..
@@
In addition, in the first part of this paper that concerns $\sf{Sealing}$, we show that several related corollaries obtained by Woodin using the stationary tower forcing, follow from Theorem 1.10 (or its proof). First, we observe that a form of $\mathsf{Sealing}$ for derived models follows from $\mathsf{Sealing}$ (see Corollary 8.1 and Corollary 8.2). Next we show that $\sf{Generically\ Correct\ Sealing}$ holds in $V[g]$ (assuming the hypothesis of Theorem 1.10), which is a generalization of $\mathsf{Sealing}$ inspired by the notion of generically correct formula (see Section 9, Definition 9.1) [^2] We do not know if $\sf{Generically\ Correct\ Sealing}$ can be established via the stationary-tower forcing.. $\sf{Generically\ Correct\ Sealing}$ implies $\mathsf{Sealing}$ (see Lemma 9.3).
@@ 1.4. The ordinal definable powerset of $\Gamma^{\infty}$
The following is one of the main theorems of this paper. First let $\mathcal{A}^{\infty}={\wp}_{uB}(\Gamma^{\infty})$. If $g$ is $V$-generic then we let $\mathcal{A}^{\infty}_{g}=(\mathcal{A}^{\infty})^{V[g]}$. If $\sf{Sealing}$ holds and $g$, $g^{\prime}$ are two consecutive generics [^7] I.e. $g^{\prime}$ is $V[g]$-generic., then we let
@@ 2. Preliminaries, notation and conventions
Theorem 2.3.
Suppose there is a class of Woodin cardinals. Then for each $\kappa$, $\Gamma^{\infty}=\operatorname{Hom}_{\kappa}$.
@@
Suppose $\mu$ is a $\eta$-complete measure and $g$ is a $<\eta$-generic. Then $\mu_{g}\in V[g]$ is the ultrafilter generated by $\mu$ in $V[g]$. Similarly, if $W$ is a set of $\eta$-complete measures and $g$ is a $<\eta$-generic then $W_{g}=\{\mu_{g}:\mu\in W\}$, and $\operatorname{TW}_{W,g}=\operatorname{TW}_{W_{g}}$. Continuing with $W$ and $g$, if $\bar{\mu}=(\mu_{s}\mid s\in\omega^{<\omega})$ is a homogeneity system then the expression $\bar{\mu}\subseteq W$ means that for every $s\in\omega^{<\omega}$, $\mu_{s}\in W$. If $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})\subseteq W$ is a homogeneity system then $\bar{\mu}_{g}=((\mu_{s})_{g}:s\in\omega^{<\omega})=_{def}(\mu_{s,g}:s\in\omega^{<\omega})$ and $S_{\mu,g}=_{def}(S_{\mu_{g}})^{V[g]}=_{def}(S_{\mu})^{V[g]}$.
@@
    $f\colon\operatorname{TW}_{Y}\rightarrow\operatorname{TW}_{R}$ is a $1$-to-$1$ Lipschitz function [^13] Recall that a function $f\colon\operatorname{TW}_{Y}\rightarrow\operatorname{TW}_{R}$ is Lipschitz if the value of $f(\vec{\mu})\upharpoonright n$ is determined by $\vec{\mu}\upharpoonright n$, for all $\vec{\mu}$ and $n$.,
@@
As in [26, Lemma 2.1] we find a tree $T$ on some $W\times U$ such that ${|W|}\leq{|Y|}$and $p[T]$ is the set of well-founded towers $\bar{\mu}\in\operatorname{TW}_{Y}$ (this fact remains true in $V[G]$ for any $<\delta$-generic $G$). We now have that $T$ is $\delta^{+}$-homogenously Suslin. The construction presented in [8] and also in [14] yields the desired $(f,R)$. Notice that the construction presented in [14] (see the discussion after [14, Exercise 5.25]) produces a map $s\mapsto{\mathcal{T}}_{s}$ with domain $W^{<\omega}$ such that for every $\vec{\mu}$, $\vec{\mu}\in p[T]$ if and only if the even branch of $\cup_{i<\omega}{\mathcal{T}}_{\vec{\mu}\restriction i}$ is ill-founded (for example, see [14, Claim 5.19], here ${\mathcal{T}}_{s}$ is a finite tree on $V_{\delta}$). Converting the map $s\mapsto{\mathcal{T}}_{s}$ into homogeneity system as is done in [14] we obtain our desired $f$ and $R$. Analyzing the conversion we see that $R$ satisfies Clause 4 above. Because each ${\mathcal{T}}_{s}$ is an iteration tree on $V_{\delta}$, we get that $R\in V_{\delta}$. We can then modify the construction of ${\mathcal{T}}_{s}$ to ensure that $f$ is $1$-to-$1$. For details see [5, Remark 9.21] and [14, Definition 5.13]. ∎
@@
Suppose $p=(f,Y,R)$ is a flipping system and $g$ is a $<{\sf{comp}}(p)$-generic. Then $f_{g}:\operatorname{TW}_{Y,g}\rightarrow\operatorname{TW}_{R,g}$ is the extension of $f$ in $V[g]$ and has the property that $f_{g}(\vec{\mu}_{g})=(f(\vec{\mu}))_{g}$. Given a tower $\vec{\mu}\in\operatorname{TW}_{Y}$, we let $f_{g}(\vec{\mu})=(f(\vec{\mu}))_{g}$.
@@ 3. Preserving universally Baire sets
Let $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})$ and $\pi\,"\mu=\bar{\nu}=(\nu_{s}:s\in\omega^{<\omega})$. Let $\delta>\kappa$ be such that $\delta\in\operatorname{rng}(\pi)$ and $\delta$ tames $W$. We can then find a flipping system $p=(f,W,R)$ that flips through $\delta$ and $p\in\operatorname{rng}(\pi)$. Let $q=(f^{\prime},W^{\prime},R^{\prime})=\pi^{-1}(p)$. Notice now that if $\vec{\tau}\in(TW_{W^{\prime}})^{M[g]}$ then $\vec{\tau}$ is well-founded (in $M[g]$) if and only if $\pi\,"\vec{\tau}$ is well-founded (see the proof of Lemma 2.11).
@@
The following is a useful lemma that can be established via the same methods.
@@
Suppose $W$ is infinitely tamed above $\kappa$, $\chi$ is such that $W\in V_{\chi}$, $\pi:M\rightarrow V_{\chi}$ is an elementary embedding with $M$ countable and transitive, and $\{W,\kappa\}\in\operatorname{rng}(\pi)$. Let $g$ be $<\pi^{-1}(\kappa)$-generic over $M$ and $\bar{\mu},\bar{\nu}\in M[g]$ be a homogeneity system such that $\bar{\mu},\bar{\nu}\subseteq M$, $\pi\,"\bar{\mu}\subseteq W$ and $\pi\,"\bar{\nu}\subseteq W$. Suppose that in $M[g]$, $S_{\bar{\mu}_{g}}=S_{\bar{\nu}_{g}}$. Then $S_{\pi\,"\bar{\mu}}=S_{\pi\,"\bar{\nu}}$.
@@
Let $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})$ and $\bar{\nu}=(\nu_{s}:s\in\omega^{<\omega})$. We let $\pi\,"\bar{\mu}=(\mu^{+}_{s}:s\in\omega^{<\omega})$ and $\pi\,"\bar{\nu}=(\nu^{+}_{s}:s\in\omega^{<\omega})$. Let $\delta_{1}>\delta_{0}>\kappa$ be such that $\{\delta_{0},\delta_{1}\}\in\operatorname{rng}(\pi)$ and both $\delta_{0}$ and $\delta_{1}$ tame $W$ (and in particular they are Woodin cardinals). Let $(\eta_{0},\eta_{1})=\pi^{-1}(\delta_{0},\delta_{1})$. We can then find a flipping system $p=(f,W,R)$ that flips through $\delta_{1}$, $p\in\operatorname{rng}(\pi)$ and $\delta_{0}<{\sf{comp}}(p)$. Let $q=(f^{\prime},W^{\prime},R^{\prime})=\pi^{-1}(p)$.
Fix now $x\in\mathbb{R}$ and using Theorem 2.5 find $\pi^{\prime}:M^{\prime}\rightarrow V_{\chi}$ and $\sigma:M\rightarrow M^{\prime}$ such that $\pi=\pi^{\prime}\circ\sigma$, ${\rm cp}(\sigma)>\pi^{-1}(\kappa)$ and for some $M^{\prime}$-generic $h\subseteq\operatorname{Col}(\omega,\sigma(\eta_{0}))$, $x\in M^{\prime}[h]$. Notice now that because $p\in\operatorname{rng}(\pi^{\prime})$ if $\vec{\tau}\in(TW_{\sigma(W^{\prime})})^{M^{\prime}[g*h]}$ then $\vec{\tau}$ is well-founded (in $M^{\prime}[g*h]$) if and only if $\pi^{\prime}\,"\vec{\tau}$ is well-founded (see the proof of Lemma 2.11). Let $\sigma^{+}:M[g]\rightarrow M^{\prime}[g]$ be the liftup of $\sigma$. We now have that,
@@
Suppose $\kappa$ is an infinite cardinal, $W$ is a set of measures that is tamed above $\kappa$ and $E$ is an extender with ${\rm cp}(E)=\kappa$. We then say that $E$ is wf-preserving for $W$161616“wf” stands for “well-founded”. (in $V$) if
@@
    in $\operatorname{Ult}(V,E)$ there is a Lipschitz function
@@
We start by fixing a flipping system $(f,W,Y)$ such that $g$ is $<{\sf{comp}}(Y)$-generic and $Y\in V_{\delta}$. Now let $h^{\prime}=j\upharpoonright\operatorname{TW}_{W}$. Since $j$ is a $\lambda$-supercompactness embedding, we have $h^{\prime}\in M$. It then follows from Lemma 2.11 that $h^{\prime}$ is wf-preserving for $W$.
@@
Let $k\colon\operatorname{Ult}(V,E)\rightarrow M$ be the factor map. Then $k\upharpoonright V_{\lambda}=\operatorname{id}$ and hence
@@
$\operatorname{Ult}(V,E)\models\varphi(W,j(W)).$
If now $h\in Ult(V,E)$ is the witness to $\varphi(W,j(W))$, then $h$ is as desired. ∎

### v3

@@ Abstract.
Inspired by core model induction, we introduce the definable powerset $\mathcal{A}^{\infty}$ of $\Gamma^{\infty}$ and use our derived model representation mentioned above to show that the theory of $L(\mathcal{A}^{\infty})$ cannot be changed by forcing (see Theorem 1.19). Working in a different direction, we also show that the theory of $L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}]$, where $\mathcal{C}$ is the club filter on ${\wp}_{\omega_{1}}(\Gamma^{\infty})$, cannot be changed by forcing (see Theorem 1.30). Proving the two aforementioned results is the first step towards showing that the theory of $L(\sf{Ord}^{\omega},\Gamma^{\infty},\mathbb{R})([\mu_{\alpha}:\alpha\in\sf{Ord}])$, where $\mu_{\alpha}$ is the club filter on ${\wp}_{\omega_{1}}(\alpha)$, cannot be changed by forcing.
The first author gratefully acknowledges that this research was funded in part by the Austrian Science Fund (FWF) [10.55776/Y1498, 10.55776/I6087, 10.55776/V844]. The second author’s work is funded by the National Science Centre, Poland under the Weave-UNISONO call in the Weave program, registration number UMO-2021/03/Y/ST1/00281. For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission. Both authors would like to thank the referee for their helpful comments which improved the presentation of this paper.
@@ 1.1. Main results
Following Woodin, we write $\Gamma^{\infty}$ for the set of universally Baire sets and, if $g$ is set generic over $V$, we write $\Gamma_{g}^{\infty}=(\Gamma^{\infty})^{V[g]}$ and $\mathbb{R}_{g}=\mathbb{R}^{V[g]}$ where $\mathbb{R}$ denotes the set of reals. Given a universally Baire set $A$ and a set generic $g\subseteq\mathbb{P}$ for some poset $\mathbb{P}$, we write $A_{g}$ for the interpretation of $A$ in $V[g]$. More precisely, letting $\kappa\geq{|\mathbb{P}|}^{+}$ be a cardinal and $(S,T)$ be any $\kappa$-absolutely complementing trees with $p[S]=A$, we set $A_{g}=(p[S])^{V[g]}$. It can be easily checked using the absoluteness of well-foundedness that $A_{g}$ is independent of $(T,S)$. The main object this paper studies is the model $L(\Gamma^{\infty},\mathbb{R})$, and our aim is to find a useful derived model representation for it [^1] By a theorem of Woodin, all models of ${\sf{AD^{+}}}$ are elementary equivalent to a model of ${\sf{AD^{+}}}$ that can be obtained as a derived model. This refers to the so-called “new” derived model of the form $L(\Gamma,\mathbb{R}^{*})$ for $V(\mathbb{R}^{*})$ a $\operatorname{Col}(\omega,{<}\lambda)$-symmetric extension of $V$ for $\lambda$ a limit of Woodin cardinals in $V$ and $\Gamma$ the set of all $A\subseteq\mathbb{R}^{*}$ such that $L(A,\mathbb{R}^{*})\models\operatorname{AD}^{+}$ and $A\in V(\mathbb{R}^{*})$. The main focus in this paper is the so-called “old” derived model as defined in Section 2..
@@
In addition, in the first part of this paper that concerns $\sf{Sealing}$, we show that several related corollaries obtained by Woodin using the stationary tower forcing, follow from Theorem 1.10 (or its proof). First, we observe that a form of $\mathsf{Sealing}$ for derived models follows from $\mathsf{Sealing}$ (see Corollary 8.1 and Corollary 8.2). Next we show that $\sf{Generically\ Correct\ Sealing}$ holds in $V[g]$ (assuming the hypothesis of Theorem 1.10), which is a generalization of $\mathsf{Sealing}$ inspired by the notion of generically correct formula (see Section 9, Definition 9.1). $\sf{Generically\ Correct\ Sealing}$ implies $\mathsf{Sealing}$ (see Lemma 9.3).
@@ 1.4. The ordinal definable powerset of $\Gamma^{\infty}$
We introduce a canonical set of subsets of $\Gamma^{\infty}$, the ordinal definable powerset of $\Gamma^{\infty}$.
@@
The following is one of the main theorems of this paper. First let $\mathcal{A}^{\infty}={\wp}_{uB}(\Gamma^{\infty})$ be the uB-powerset. If $g$ is $V$-generic then we let $\mathcal{A}^{\infty}_{g}=(\mathcal{A}^{\infty})^{V[g]}$. If $\sf{Sealing}$ holds and $g$, $g^{\prime}$ are two consecutive generics [^6] I.e. $g^{\prime}$ is $V[g]$-generic., then we let
@@ 2. Preliminaries, notation and conventions
Theorem 2.3 (Martin-Steel-Woodin).
Suppose there is a class of Woodin cardinals. Then $\Gamma^{\infty}=\cap_{\kappa\in\sf{Ord}}\operatorname{Hom}_{\kappa}$.
@@
Suppose $\mu$ is an $\eta$-complete measure for some $\eta$ and $g$ is ${<}\eta$-generic. Then $\mu_{g}\in V[g]$ is the ultrafilter generated by $\mu$ in $V[g]$. Similarly, if $W$ is a set of $\eta$-complete measures and $g$ is ${<}\eta$-generic then $W_{g}=\{\mu_{g}:\mu\in W\}$, and $\operatorname{TW}_{W,g}=\operatorname{TW}_{W_{g}}$. Continuing with $W$ and $g$, if $\bar{\mu}=(\mu_{s}\mid s\in\omega^{<\omega})$ is a homogeneity system then the expression $\bar{\mu}\subseteq W$ means that for every $s\in\omega^{<\omega}$, $\mu_{s}\in W$. If $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})\subseteq W$ is a homogeneity system, then $\bar{\mu}_{g}=((\mu_{s})_{g}:s\in\omega^{<\omega})=_{def}(\mu_{s,g}:s\in\omega^{<\omega})$ and $S_{\mu,g}=_{def}(S_{\mu_{g}})^{V[g]}=_{def}(S_{\mu})^{V[g]}$.
@@
    $f\colon\operatorname{TW}_{Y}\rightarrow\operatorname{TW}_{R}$ is a $1$-to-$1$ Lipschitz function [^12] Recall that a function $f\colon\operatorname{TW}_{Y}\rightarrow\operatorname{TW}_{R}$ is Lipschitz if the value of $f(\vec{\mu})\upharpoonright n$ is determined by $\vec{\mu}\upharpoonright n$, for all $\vec{\mu}$ and $n$., and
@@
As in [26, Lemma 2.1] we find a tree $T$ on some $W\times U$ such that ${|W|}\leq{|Y|}$ and $p[T]$ is the set of well-founded towers $\vec{\mu}\in\operatorname{TW}_{Y}$ (this fact remains true in $V[G]$ for any ${<}\delta$-generic $G$). We now have that $T$ is $\delta^{+}$-homogenously Suslin. The construction presented in [8] and also in [14] yields the desired $(f,R)$. Notice that the construction presented in [14] (see the discussion after [14, Exercise 5.25]) produces a map $s\mapsto{\mathcal{T}}_{s}$ with domain $W^{<\omega}$ such that for every $\vec{\mu}$, $\vec{\mu}\in p[T]$ if and only if the even branch of $\bigcup_{i<\omega}{\mathcal{T}}_{\vec{\mu}\restriction i}$ is ill-founded (for example, see [14, Claim 5.19], here ${\mathcal{T}}_{s}$ is a finite tree on $V_{\delta}$). Converting the map $s\mapsto{\mathcal{T}}_{s}$ into a homogeneity system as it is done in [14] we obtain our desired $f$ and $R$. Analyzing the conversion we see that $R$ satisfies Clause (4) above. Because each ${\mathcal{T}}_{s}$ is an iteration tree on $V_{\delta}$, we get that $R\in V_{\delta}$. We can then modify the construction of ${\mathcal{T}}_{s}$ to ensure that $f$ is $1$-to-$1$. For details see [6, Remark 9.21] and [14, Definition 5.13]. ∎
@@
Suppose $p=(f,Y,R)$ is a flipping system and $g$ is ${<}{\sf{comp}}(p)$-generic. Then $f_{g}:\operatorname{TW}_{Y,g}\rightarrow\operatorname{TW}_{R,g}$ is the extension of $f$ to $V[g]$ and has the property that $f_{g}(\vec{\mu}_{g})=(f(\vec{\mu}))_{g}$. Given a tower $\vec{\mu}\in\operatorname{TW}_{Y}$, we let $f_{g}(\vec{\mu})=(f(\vec{\mu}))_{g}$.
@@ 3. Preserving universally Baire sets
Let $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})$ and $\pi\,"\bar{\mu}=\bar{\nu}=(\nu_{s}:s\in\omega^{<\omega})$. Let $\delta>\kappa$ be such that $\delta\in\operatorname{rng}(\pi)$ and $\delta$ tames $W$. We can then find a flipping system $p=(f,W,R)$ that flips through $\delta$ with $p\in\operatorname{rng}(\pi)$. Let $q=(f^{\prime},W^{\prime},R^{\prime})=\pi^{-1}(p)$. Notice now that if $\vec{\tau}\in(TW_{W^{\prime}})^{M[g]}$ then $\vec{\tau}$ is well-founded (in $M[g]$) if and only if $\pi\,"\vec{\tau}$ is well-founded (see the proof of Lemma 2.11).
@@
The following two lemmas are useful lemmas that can be established using the same methods.
@@
Suppose $W$ is a set of measures and $\delta$ tames $W$. Suppose $g$ is ${<}\delta$-generic, and $\bar{\mu}$, $\bar{\nu}$ in $V[g]$ are two homogeneity systems such that $\bar{\mu}\subseteq W$ and $\bar{\nu}\subseteq W$. Let $h$ be ${<}\delta$-generic over $V[g]$. Then $V[g]\models S_{\bar{\mu}_{g}}=S_{\bar{\nu}_{g}}$ if and only if $V[g*h]\models S_{\bar{\mu}_{g}}=S_{\bar{\nu}_{g}}$.
@@
Toward a contradiction, suppose not. Let $\chi$ be a limit cardinal such that $W\in V_{\chi}$. Let $\pi:M\rightarrow V_{\chi}[g]$ be an elementary embedding with $M$ countable and transitive, and $\{W,\delta\}\in\operatorname{rng}(\pi)$. Let $(\bar{W},\bar{g},\bar{\delta},\bar{\tau},\bar{\xi})=\pi^{-1}(W,g,\delta,\bar{\mu},\bar{\nu})$. Let now $k$ be $M$-generic for a poset of size $<\bar{\delta}$, and suppose that $M[k]\models S_{\bar{\tau}_{\bar{g}*k}}\not=S_{\bar{\xi}_{\bar{g}*k}}$. Suppose that there is $x\in M[k]\cap\mathbb{R}_{g}$ such that $x\in S_{\bar{\tau}_{\bar{g}*k}}$ but $x\not\in S_{\bar{\xi}_{\bar{g}*k}}$. Then it follows from Lemma 3.2 that $x\in S_{\bar{\mu}_{g}}$ but $x\not\in S_{\bar{\nu}_{g}}$, which is a contradiction. ∎
Lemma 3.4.
Suppose $W$ is infinitely tamed above $\kappa$, $\chi$ is such that $W\in V_{\chi}$, $\pi:M\rightarrow V_{\chi}$ is an elementary embedding with $M$ countable and transitive, and $\{W,\kappa\}\in\operatorname{rng}(\pi)$. Let $g$ be ${<}\pi^{-1}(\kappa)$-generic over $M$ and $\bar{\mu},\bar{\nu}\in M[g]$ be homogeneity systems such that $\bar{\mu},\bar{\nu}\subseteq M$, $\pi\,"\bar{\mu}\subseteq W$ and $\pi\,"\bar{\nu}\subseteq W$. Suppose that in $M[g]$, $S_{\bar{\mu}_{g}}=S_{\bar{\nu}_{g}}$. Then $S_{\pi\,"\bar{\mu}}=S_{\pi\,"\bar{\nu}}$.
Proof.
Let $\bar{\mu}=(\mu_{s}:s\in\omega^{<\omega})$ and $\bar{\nu}=(\nu_{s}:s\in\omega^{<\omega})$. We let $\pi\,"\bar{\mu}=(\mu^{+}_{s}:s\in\omega^{<\omega})$ and $\pi\,"\bar{\nu}=(\nu^{+}_{s}:s\in\omega^{<\omega})$. Let $\delta_{1}>\delta_{0}>\kappa$ be such that $\{\delta_{0},\delta_{1}\}\in\operatorname{rng}(\pi)$ and both $\delta_{0}$ and $\delta_{1}$ tame $W$ (and in particular they are Woodin cardinals). Let $(\eta_{0},\eta_{1})=\pi^{-1}(\delta_{0},\delta_{1})$. We can then find a flipping system $p=(f,W,R)$ that flips through $\delta_{1}$ with $p\in\operatorname{rng}(\pi)$ and $\delta_{0}<{\sf{comp}}(p)$. Let $q=(f^{\prime},W^{\prime},R^{\prime})=\pi^{-1}(p)$.
Fix now $x\in\mathbb{R}$ and using Theorem 2.5 find $\pi^{\prime}:M^{\prime}\rightarrow V_{\chi}$ and $\sigma:M\rightarrow M^{\prime}$ such that $\pi=\pi^{\prime}\circ\sigma$, ${\rm cp}(\sigma)>\pi^{-1}(\kappa)>{|g|}$ and for some $M^{\prime}[g]$-generic $h\subseteq\operatorname{Col}(\omega,\sigma(\eta_{0}))$, $x\in M^{\prime}[g*h]$. Notice now that because $p\in\operatorname{rng}(\pi^{\prime})$ if $\vec{\tau}\in(TW_{\sigma(W^{\prime})})^{M^{\prime}[g*h]}$ then $\vec{\tau}$ is well-founded (in $M^{\prime}[g*h]$) if and only if $\pi^{\prime}\,"\vec{\tau}$ is well-founded (see the proof of Lemma 2.11). Let $\sigma^{+}:M[g]\rightarrow M^{\prime}[g]$ be the liftup of $\sigma$. We now have that,
@@
Suppose $\kappa$ is an infinite cardinal, $W$ is a set of measures that is tamed above $\kappa$ and $E$ is an extender with ${\rm cp}(E)=\kappa$. We then say that $E$ is wf-preserving for $W$151515“wf” stands for “well-founded”. (in $V$) if
@@
    in $Ult(V,E)$ there is a Lipschitz function
@@
We start by appealing to Lemma 2.9 and fixing a flipping system $(f,W,Y)$ such that $g$ is $<{\sf{comp}}(Y)$-generic and $Y\in V_{\delta}$. Now let $h^{\prime}=j\upharpoonright\operatorname{TW}_{W}$. Since $j$ is a $\lambda$-supercompactness embedding, we have $h^{\prime}\in M$. It then follows from Lemma 2.11 that $h^{\prime}$ is wf-preserving for $W$.
@@
Let $k\colon Ult(V,E)\rightarrow M$ be the factor map. Then $k\upharpoonright V_{\lambda}=\operatorname{id}$ and hence
@@
$Ult(V,E)\models\varphi(W,\pi_{E}(W)).$
If now $h\in Ult(V,E)$ is the witness to $\varphi(W,\pi_{E}(W))$, then $h$ is as desired. ∎